If (x-2)(to the second power)= 25 and x is less than 0, what is the value of x.
Answer by mananth(16946)

(x-2)^2=25
take the square root
x-2 = +/- 5
x= 2+5 OR x=2-5
x=7 OR -3
x<0
so x=-3